OpenStudy (anonymous):

I have a critcal value, which i dtermined is a max. Now i want to plug it back into the original function inorder to get the y value. The orginal function is (x/2)+cosx. The critical point is pi/4.The books gives me an answer of root 2. How did they get this?

OpenStudy (anonymous):

hrmm. let me do the calc on it.

OpenStudy (anonymous):

derivitive of (x/2) is 1/2 and cosx is -sinx so 1/2-sinx=0 sinx = 1/2 at pi/6 ? been awhile am i doing that right? take derivitive set to 0 to find critical points?

OpenStudy (anonymous):

if so then pi/6 is max, pi/4 was slightly off of zero slope

OpenStudy (anonymous):

still dont get root 2 though

OpenStudy (anonymous):

Maybe the book is wrong...?

OpenStudy (anonymous):

yeah, that the problem, i cant do the computation, i dont know how to deal with this pi/4 term

OpenStudy (anonymous):

are you looking for a global max?

OpenStudy (anonymous):

Sin pi/4 = 1/root 2

OpenStudy (anonymous):

no , i know that the critical point gives me a rel. max. and now i want to plug the critical point bakc into the orginal equation inorder to get the y value

OpenStudy (anonymous):

i get pi/6 if i just solve 1/2=sinx so critical point is pi/6

OpenStudy (anonymous):

and y = 1.1278

OpenStudy (anonymous):

something like this: \[((\pi/6)/2)+\cos(\pi/6)\] No, your right the critical point is pi/6, thats what i meant to write, but now i want to find the y value, and to to that i need to plug pi/6 into the orginal equation, which gives me the equation above

OpenStudy (anonymous):

The books says the answer should be (pi+6root3)/12)

OpenStudy (anonymous):

your book must has to be wrong i guess?

OpenStudy (anonymous):

it's asking for the point of the relative max?

OpenStudy (anonymous):

I got it now, thanks

OpenStudy (anonymous):

oh what was the problem?

OpenStudy (anonymous):

laziness

OpenStudy (anonymous):

lol... k